# Quantum Signatures of Chaos from Free Probability

### v1

@@ 3.2.1 Free additive convolution for arbitrary probability measures
In this section, we briefly review the method described in Section 5 of speicher2019lecturenotesfreeprobability to compute the free additive convolution of arbitrary probability measures. This approach is useful for studying more general algebras, beyond those involving spin-1/2 and spin-1 operators that appear in the models considered in this work.
@@
From this the eigenvalue distribution of the sum of these two operators can be obtained via the inverse Stieltjes transform:
@@
These functions satisfy the following fixed-point equation for $\omega_{a}(z)\in\mathbb{C}^{+}$ (the upper half-plane) speicher2019lecturenotesfreeprobability :
@@
This method only requires knowledge of the Cauchy transform of $a$ and $b$, which is not too difficult to obtain. In Section 3.5, we employ this procedure to compute the free additive convolution for sums of operators with spin-$3/2$, $2$, and $5/2$.
3.3 Asymptotic freeness and the vanishing of OTOCs
Under these conditions, we expect that for two operators $A,B\in\mathcal{A}$, if the dynamics is chaotic, the time-evolved operator $B(t)$ becomes asymptotically free from $A(0)$ with respect to the state (46) at sufficiently large times. This expectation is motivated by the fact that in a chaotic system, the time evolution operator $e^{-iHt}$ is well-approximated by a Haar-random unitary, i.e.,
$e^{-iHt}\approx U_{\text{Haar}}.$
3.4 Incorporating Freeness into the Quantum Ergodic Hierarchy
F-systems.
A quantum dynamical system $(\mathcal{A},\varphi,T_{t})$ is said to be an F-system or to possess the F-property if, for any operators $A,B\in\mathcal{A}$, the operators $A$ and $B(t)$ are free in the limit $t\to\infty$. We also refer to the F-property as freeness.
It is clear that the F-property implies quantum mixing, but it is unclear whether this condition is stronger or weaker than quantum $K$-mixing. We suspect that the F-property, or freeness, lies between quantum mixing and quantum $K$-mixing in the ergodic hierarchy. See Figure 6.
3.5 Operator statistics as a signature of asymptotic freeness
The fact that freeness can only emerge in the large $\mathcal{D}$ [^13] We remind the reader that $\mathcal{D}$ denotes the rank of the operators in the algebra. limit can be understood from an algebraic perspective, particularly in the context of the quantum ergodic hierarchy. First, observe that freeness implies 2-mixing. Since mixing is a property that arises only for type III algebras, it follows that freeness can occur only for type III algebras. For finite $\mathcal{D}$, however, the algebra of observables is always of type I, and therefore freeness cannot occur. As a result, mixed free cumulants of centered variables ($A-\langle A\rangle$ and $B(t)-\langle B(t)\rangle$) will not generically vanish for finite $\mathcal{D}$. However, if the dynamics is chaotic, these cumulants are expected to approach a small residual value that scales inversely with the number of degrees of freedom in the system. This expectation was indeed shown to be true for OTOCs in Huang:2017fng . Furthermore, if two operators $A$ and $B(t)$ are free, the eigenvalue density of their sum $\rho_{A(0)+B(t)}$ is expected to follow the free convolution prediction, which can emerge even for finite, but sufficiently large $\mathcal{D}$, and serves as a “smoking gun” of asymptotic freeness Chen:2024zfj . These concepts are illustrated schematically in Figure 6.
Figure 6: Illustration of the quantum and classical ergodic hierarchies, with freeness (F-property) positioned between quantum mixing and K-mixing. This placement is not fully established and remains an open question, as the precise relationship between these concepts is still under investigation. $N$ denotes the number of degrees of freedom of the system. The corresponding von Neumann algebraic structures associated with different degrees of quantum chaos are indicated. Note that the quantum hierarchy is only meaningful in the $1/N\rightarrow 0$ limit, as for any finite $N$, the algebra remains type I, preventing even ergodicity. Nonetheless, a finite-$N$ system can exhibit signatures of its asymptotic behavior. For example, the emergence of operator spectral statistics governed by free probability theory serves as a robust indicator of asymptotic freeness.
We illustrate this idea by studying the spectrum of operators of the form
$X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$ (47)
@@ 3.5 Operator statistics as a signature of asymptotic freeness
where the time evolution operator is $U=e^{-it\,M}$, with $M$ being a random matrix drawn from a Gaussian orthogonal ensemble (GOE), and $X^{(s)}_{i}$ are generalized spin-$s$ operators (see Sec. 5 for a precise definition). For a chain with $L$ sites, these operators have dimension $d_{s}=(2s+1)^{L}$. The variance of $M$ was fixed at $2/d_{s}$. Figure 7 shows the spectrum of $X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$ for spin-$1/2$ operators (left panel) and spin-1 operators (right panel), along with the free convolution prediction for the corresponding operators (see eqs. (33) and (39)). Figure 8 shows the spectrum of $X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$ for spin-$3/2$, spin-2, and spin-$5/2$ operators, along with the numerical free convolution prediction for the corresponding operators, obtained using the method described in Section 3.2.1. These results demonstrate that when the dynamics are driven by a random matrix, the spectrum of $X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$ aligns with the predictions of free probability. This is expected in light of (20).
@@
(c) spin-5/2
@@
Figure 8: Density of eigenvalues for the operator sum $X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$, where $X^{(s)}$ spin-$s$ operators and $U=e^{-itM}$ represents the time evolution operator. Here we set $t=50$. The matrix $M$ is drawn from a GOE with zero mean and variance $2/d_{s}$, where $d_{s}=(2s+1)^{L}$ is the Hilbert space dimension for a spin-$s$ chain with $L$ sites. The histograms are based on 1000 independent realizations of $M$. Panels correspond to (a) spin-3/2 operators with $L=5$, (b) spin-2 operators with $L=4$, and (c) spin-5/2 operators with $L=4$. The operators are normalized such that their distribution reads $\rho_{s=3/2}=\frac{1}{4}(\delta_{-3}+\delta_{-1}+\delta_{1}+\delta_{3})$, $\rho_{s=2}=\frac{1}{5}(\delta_{-2}+\delta_{-1}+\delta_{0}+\delta_{1}+\delta_{2})$, and $\rho_{s=5/2}=\frac{1}{6}(\delta_{-5}+\delta_{-3}+\delta_{-1}+\delta_{1}+\delta_{3}+\delta_{5})$. The blue curves represent the free probability prediction obtained using the method described in Section 3.2.1.
Next three sections describe how one can use the free additive convolution prediction to probe chaos in quantum many-body systems.
@@ 4.1 Spectral statistics of the Hamiltonian
[…]
@@ 4.2 Approximate emergence of asymptotic freeness
In this section, we examine the spectral statistics of operators of the form $Z_{i}(0)+Z_{j}(t)$, where the generalized Pauli operator is time-evolved under the Hamiltonian (53). Both $Z_{i}(0)$ and $Z_{j}(t)$ have eigenvalues ±$\pm 1$, characterized by a Bernoulli distribution. When these operators are asymptotically free, the spectrum of their sum is expected to follow an arcsine distribution (see Example 1 in section 3.2):
$f(\lambda)=\begin{cases}\frac{1}{\pi\sqrt{4-\lambda^{2}}},&\text{if }0<\lambda<1,\\ 0,&\text{otherwise.}\end{cases}$ (54)
@@ 4.3 Fluctuations on top of the arcsine law
Figure 16: Eigenvalue density, level spacing distribution, and $r$-parameter distribution for the operator $X_{1}+X_{10}(t)$ based on 1000 realizations of the Hamiltonian (53) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $L=10$ and $t=50$. The $r$-parameter distribution was fitted using the form (83), yielding the Wigner parameter $\nu=2.000\pm 0.002$. The solid lines represent (a) the arcsine distribution, (b) the Wigner-Dyson distribution with $\nu=2$, and (c) the Wigner $r$-parameter distribution with $\nu=2$. The corresponding average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% at each edge.
@@
Figure 17: Eigenvalue density, level spacing distribution, and $r$-parameter distribution for the operator $Z_{1}+Z_{10}(t)$ based on 1000 realizations of the Hamiltonian (53) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $L=10$ and $t=50$. The $r$-parameter distribution was fitted using the form (83), yielding the Wigner parameter $\nu=2.000\pm 0.002$. The solid lines represent (a) the arcsine distribution, (b) the Wigner-Dyson distribution with $\nu=2$, and (c) the Wigner $r$-parameter distribution with $\nu=2$. The corresponding average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% at each edge.
@@
Figure 18: Eigenvalue density, level spacing distribution, and $r$-parameter distribution for the operator $X_{1}+X_{10}(t)$ based on 1000 realizations of the Hamiltonian (53) at the near-integrable point $(h_{x},h_{z},g)=(-1,0.005,0.2)$, with $L=10$ and $t=5\times 10^{4}$. The $r$-parameter distribution was fitted using the form (83), yielding the Wigner parameter $\nu=2.001\pm 0.002$. The solid lines represent (a) the arcsine distribution, (b) the Wigner-Dyson distribution with $\nu=2$, and (c) the Wigner $r$-parameter distribution with $\nu=2$. The corresponding average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% at each edge.
@@
Figure 19: Eigenvalue density, level spacing distribution, and $r$-parameter distribution for the operator $Z_{1}+Z_{10}(t)$ based on 1000 realizations of the Hamiltonian (53) at the near-integrable point $(h_{x},h_{z},g)=(-1,0.005,0.2)$, with $L=10$ and $t=5\times 10^{4}$. The $r$-parameter distribution was fitted using the form (83), yielding the Wigner parameter $\nu=2.001\pm 0.002$. The solid lines represent (a) the arcsine distribution, (b) the Wigner-Dyson distribution with $\nu=2$, and (c) the Wigner $r$-parameter distribution with $\nu=2$. The corresponding average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% at each edge.
@@
In finite numerical simulations, free cumulants in the model (53) do not vanish due to finite-size effects. The above results show that while the dynamics drives $Z_{1}$ and $Z_{10}(t)$ towards asymptotic freeness – producing a spectrum for their sum that respects the free probability prediction – the same is not entirely true for $X$-Pauli operators, in which one observes small deviations from the free probability prediction even at the chaotic point. This suggests that the finite size effects for mixed cumulants involving $X$-Pauli operators are larger than the ones involving $Z$-Pauli operator, which is indeed what is observed in numerical simulations.

### v3

@@ 4.1.1 Free additive convolution for arbitrary probability measures
In this section, we briefly review the method described in Section 5 of speicher2019lecture to compute the free additive convolution of arbitrary probability measures. See also Appendix B of Pollock:2025acf . This approach is useful for studying more general algebras, beyond those involving spin-1/2 and spin-1 operators that appear in the models considered in this work.
@@
From this, the eigenvalue distribution of the sum of these two operators can be obtained via the inverse Stieltjes transformation formula:
@@
These functions satisfy the following fixed-point equation for $\omega_{a}(z)\in\mathbb{C}^{+}$ (the upper half-plane) speicher2019lecture :
@@
This method only requires knowledge of the Cauchy transform of $a$ and $b$, which is not too difficult to obtain. Below, we present a couple of detailed examples illustrating how to apply the subordination function method to numerically compute the free convolution prediction.
Example 3. Sum of Spin-2 operators
To exemplify the use of subordination function methods, we now consider the sum of two free spin-2 operators, which one with eigenvalue distribution given by
$\rho_{s=2}(\lambda)=\frac{\delta(\lambda-2)+\delta(\lambda-1)+\delta(\lambda)+\delta(\lambda+1)+\delta(\lambda+2)}{5}\,.$ (60)
The corresponding Cauchy distribution is given by
$G_{s=2}(z)=\int d\lambda\frac{\rho_{s=2}(\lambda)}{z-\lambda}=\frac{5\left(z^{2}-3\right)z^{2}+4}{5\left(z^{5}-5z^{3}+4z\right)}\,.$ (61)
The auxiliary function can then be obtained as
$H_{s=2}(z)=\frac{1}{G_{s=2}(z)}-z=\frac{2z\left(8-5z^{2}\right)}{5\left(z^{2}-3\right)z^{2}+4}\,.$ (62)
Now, we can numerically solve (59) for $H_{a}(z)$ given by (62) to find $\omega_{a}(z)$. Then we can employ Eqs. (56) and (57) to numerically find the distribution for the sum of two free spin-2 operators.
Example 4. Sum of Spin-$s$ operators in the large $s$ limit
We now consider the sum of two free spin-$s$ operators, denoted as $X^{(s)}$, in the large $s$ limit. The density of eigenvalues of each operator is given by:
$\rho_{X^{(s)}}(\lambda)=\frac{1}{2s+1}\sum_{j=-s}^{s}\delta(\lambda-j)\,.$ (63)
To simplify the analysis, it is convenient to consider the normalized operator $x^{(s)}=X^{(s)}/s$, whose density of eigenvalues in the large-$s$ limit can be well-approximated by a uniform distribution with support between -1 and 1, namely:
$\rho_{x^{(s)}}(\lambda)=\begin{cases}\frac{1}{2},~~~\lambda\in[-1,1]\,,\\ 0,~~~\text{otherwise}\,.\end{cases}$ (64)
The Cauchy transformation corresponding to this distribution is simply given by
$G_{\rho^{(s)}}(z)=\frac{1}{2}\log\left(\frac{z+1}{z-1}\right)\,.$ (65)
From this, we get the expression for the $R$-transformation to be
$R_{\rho^{(s)}}(z)=\coth z-\frac{1}{z}~.$ (66)
Now considering the free additive convolution of two spin-$s$ operators having the same large value of $s$, $x_{3}^{(s)}=x_{1}^{(s)}+x_{2}^{(s)}$ the $R$-transformation of the sum is given by $R_{\rho_{3}^{(s)}}(z)=2R_{\rho_{1}^{(s)}}(z)$. Therefore, the Cauchy transformation ($G_{\rho_{3}^{(s)}}(z)$) of the convolved distribution $\rho_{3}(\lambda)=\rho_{x_{1}^{(s)}}(\lambda)\boxplus\rho_{x_{2}^{(s)}}(\lambda)$ satisfies the following transcendental equation Potters ; speicher2009free , -p0s
$\mathcal{B}_{\rho_{3}^{(s)}}(G_{\rho_{3}^{(s)}}(z))=z~~\rightarrow~~2\coth(G_{\rho_{3}^{(s)}}(z))-\frac{1}{G_{\rho_{3}^{(s)}}(z)}-z=0~.$ (67)
This equation can not be solved exactly for the Cauchy transformation. However, one can determine the range of the spectrum of the operator $x_{3}^{(s)}$ by plotting the inverse of $G_{\rho_{3}}(z)$, which has a local maximum and a local minimum. The eigenvalues of $x_{3}^{(s)}$ lie within this region where there is no real solution of the equation $\mathcal{B}_{\rho_{3}^{(s)}}(G_{\rho_{3}^{(s)}}(z))=z$. As can be seen from the plot in Fig. 8, the edges of the spectrum are approximately at 1.543-1.543 and $1.543$.
Figure 8: Plot of the function $\mathcal{B}_{\rho_{3}^{(s)}}(G_{\rho_{3}^{(s)}}(z))(=z)$. The edges of the spectrum of $x_{3}^{(s)}$ are located between a local maximum and minimum of this function, shown by the dashed black lines. We have shown the minimum of the function more clearly in the inset.
To obtain the density of eigenvalues, we use the subordination method described in section 4.1.1. The auxiliary function (58) corresponding to the Cauchy transformation in (65) reads
$H_{x^{(s)}}(z)=\frac{2-z\log\left(\frac{z+1}{z-1}\right)}{\log\left(\frac{z+1}{z-1}\right)}\,.$ (68)
The same approximation in terms of a uniform distribution works well for the non-normalized operator $X^{(s)}$, for which the corresponding auxiliary function is given by
$H_{X^{(s)}}(z)=\frac{2s+1-z\log\left(\frac{z+s}{z-s}\right)}{\log\left(\frac{z+s}{z-s}\right)}\,.$ (69)
Again, we can employ Eqs. (56) and (57) to numerically find the distribution for the sum of two free spin-$s$ operators in the large $s$ limit.
Numerics versus Free probability Predictions
Now, we apply the subordination function methods explained above to compute the free additive convolution for the sums of spin-$s$ operators for several values of $s$. We compare the free probability prediction with numerical results obtained by direct diagonalization of two operators of the form
$X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U,$ (70)
@@
where the time-evolution operator is $U=e^{-it\,M}$, with $M$ drawn from a Gaussian orthogonal ensemble (GOE), and $X^{(s)}_{i}$ are generalized spin-$s$ operators (see Section 6 for a precise definition). We compare the numerically obtained spectra with the free additive convolution prediction discussed in Section 4.1.1. For a chain with $L$ sites, these operators have dimension $d_{s}=(2s+1)^{L}$. The variance of $M$ was fixed at $2/d_{s}$. Figure 9 shows the spectrum of $X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$ for spin-$1/2$ operators (left panel) and spin-1 operators (right panel), along with the free convolution prediction for the corresponding operators (see eqs. (47) and (53)). Figure 10 shows the spectrum of $X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$ for spin-$3/2$, spin-2, spin-$5/2$ and spin-100 operators, along with the numerical free convolution prediction for the corresponding operators, obtained using the method described in Section 4.1.1. These results demonstrate that when the dynamics are driven by a random matrix, the spectrum of $X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$ aligns with the predictions of free probability. This is expected in light of (21) and the discussion presented in section 3.4, but has not yet been formally proved.
@@
Refer to caption Refer to caption
(c) spin-5/2 (d) spin-100
@@
Figure 10: Density of eigenvalues for the operator sum $X^{(s)}_{i}+U^{\dagger}X^{(s)}_{i}U$, where $X^{(s)}$ are spin-$s$ operators and $U=e^{-itM}$ represents the time evolution operator. Here we set $t=50$. The matrix $M$ is drawn from a GOE with zero mean and variance $2/d_{s}$, where $d_{s}=(2s+1)^{L}$ is the Hilbert space dimension for a spin-$s$ chain with $L$ sites. Panels correspond to (a) spin-3/2 operators with $L=5$, (b) spin-2 operators with $L=4$, (c) spin-5/2 operators with $L=4$, and (d) spin-100 operators with $L=1$. The histograms in (a), (b), and (c) are based on 1000 independent realizations of $M$, while (d) is based on 5000 realizations. The operators are normalized such that their distributions read: $\rho_{s=3/2}=\frac{1}{4}(\delta_{-3}+\delta_{-1}+\delta_{1}+\delta_{3})$, $\rho_{s=2}=\frac{1}{5}(\delta_{-2}+\delta_{-1}+\delta_{0}+\delta_{1}+\delta_{2})$, $\rho_{s=5/2}=\frac{1}{6}(\delta_{-5}+\delta_{-3}+\delta_{-1}+\delta_{1}+\delta_{3}+\delta_{5})$, and $\rho_{X^{(s)}}(\lambda)=\frac{1}{2s+1}$ for $\lambda\in[-s,s]$, and zero otherwise. The blue curves represent the free probability prediction obtained using the subordination function method.
4.2 Decorrelated ensembles and the role of eigenvectors
When considering any of the Hermitian Gaussian random matrix ensembles, there are two key aspects that these share: correlations between their eigenvalues, which result in the Wigner–Dyson level-spacing distribution and sine kernel universality arising from energy-density pair correlations, and delocalization of their eigenvectors. Eigenvectors of Gaussian RMTs are uniformly distributed over the corresponding orthogonal/unitary/ symplectic group according to the Haar measure and become isotropically delocalized with respect to the standard basis in $\mathbb{R}^{{\mathcal{D}}}$ Erdos_2009 ; Orourke:2016eig . In contrast, the eigenvectors of the Poissonian ensemble are completely localized in the standard basis in $\mathbb{R}^{{\mathcal{D}}}$. These two aspects are intimately tied to each other and form part of the same underlying RMT behaviour Denton:2019pka . It is thus difficult to imagine a situation in which we could have one of these properties without the other. However, one can still ask whether both aspects are necessary for triggering the asymptotic freeness between operators when one of them evolves in time via a Hamiltonian drawn from such Gaussian RMT ensembles. One way to test this, for example, is to artificially remove correlations from the eigenvalues while retaining the delocalized properties of the eigenvectors as well as the same average density of states, which in this case in the large-$\mathcal{D}$} limit is given by the Wigner semicircle.
To be precise, suppose that we draw one Hamiltonian $H$ from a Hermitian Gaussian RM ensemble (with fixed variance $\sigma^{2}_{\textrm{GUE}}=1/{\mathcal{D}}$, $\sigma^{2}_{\textrm{GOE}}=2/{\mathcal{D}}$, $\sigma^{2}_{\textrm{GSE}}=1/(2{\mathcal{D}})$ and zero mean) and find its eigenvectors $|n\rangle$ and eigenvalues $E_{n}$. For simplicity, we can imagine that these eigenvalues are arranged in an ascending order with respect to $n$ together with their corresponding eigenvectors. The eigenvalues of $H$ will give rise to a discrete density of states $\rho(E)$ whose average $\overline{\rho(E)}$ will tend to the Wigner semicircle law $\rho_{W}(E)=\sqrt{4-E^{2}}/(2\pi)$ (for $|E|\leq 2$) in the thermodynamic limit $\mathcal{D}$$\rightarrow\infty$. The true eigenvalues $E_{n}$ of $H$ will show correlations in the way described above due to fluctuations around the average distribution $\overline{\rho(E)}$. However, if we now independently draw $\mathcal{D}$} variables $\{E_{n}^{\prime}\}$ from $\overline{\rho(E)}$ (or equivalently from the Wigner semicircle distribution $\rho_{W}(E)$) then these will no longer show the above correlations, while still displaying the same average density distribution as the original correlated eigenvalues. Because of this we can call them “decorrelated” eigenvalues. In particular, the level-spacing distribution will be Poissonian. We can now construct a new Hamiltonian $H^{\prime}$ from the original eigenvectors $|n\rangle$ and the new set of decorrelated eigenvalues $E^{\prime}_{n}$ arranged ascendingly. Moreover, we can perform this procedure for as many Hamiltonians as required, since all Hamiltonians drawn from Gaussian RMTs are diagonalizable and have real eigenvalues. In the context of this work, we will refer to such a class of random matrix ensembles, “decorrelated” Gaussian ensembles. [^21] This is same spirit as the “Poissonian Hamiltonian ensembles” described in Balasubramanian:2022tpr . We now ask whether time evolution under such a class of Hamiltonians produces asymptotic freeness or not.
In Appendix H we numerically show that at least in the case of decorrelated GOE matrices, and for operators of spins $1/2$ and $1$, the asymptotic freeness prediction arises for operator sums of the form $X_{i}(0)+X_{j}(t)$ for $i=j$ and $i\neq j$. Moreover, the operator statistics resulting from the time evolution generated by the decorrelated GOE matrices matches the statistics obtained from the true time evolution by the GOE matrices. This seems to suggest that, at least in the context of GOE matrices, time evolution under Hamiltonians with delocalized eigenvectors is sufficient to give rise to asymptotic freeness.
This analysis, along with the emergence of free convolution near integrability for certain operators, suggests that freeness can arise even without random matrix behavior, depending on the choice of operators. Since freeness reflects strong operator growth and scrambling – in the sense of vanishing of OTOCs and their higher order counterparts – this highlights a subtle distinction between operator growth and chaos defined via random matrix universality. While random matrix behavior ensures freeness for any operators (the F-property), a weaker, operator-dependent form of freeness may still emerge in its absence. This aligns with the findings of Dowling:2023hqc , which show that scrambling is necessary but not sufficient for random matrix chaos. From the perspective of freeness, random matrix chaos implies universal freeness, but freeness itself can appear in a more limited, operator-specific way when conserved quantities obstruct full random matrix universality.
@@ 5.1 Spectral statistics of the Hamiltonian
Refer to caption $t=1$ Refer to captiont=5t=5
Refer to captiont=7t=7 Refer to captiont=20t=20
@@ 5.2 Approximate emergence of asymptotic freeness
In this section, we examine the spectral statistics of operators of the form $Z_{i}(0)+Z_{j}(t)$, where the generalized Pauli operator is time-evolved under the Hamiltonian (76). Both $Z_{i}(0)$ and $Z_{j}(t)$ have eigenvalues $\pm 1$, characterized by a Bernoulli distribution. When these operators are asymptotically free, the spectrum of their sum is expected to follow an arcsine distribution (see Example 1 in section 4.1):
$f(\lambda)=\begin{cases}\frac{1}{\pi\sqrt{4-\lambda^{2}}},&\text{if }-2<\lambda<2,\\ 0,&\text{otherwise.}\end{cases}$ (77)
@@ 5.3 Fluctuations on top of the arcsine law
Figure 18: Eigenvalue density, level spacing distribution, and $r$-parameter distribution for the operator $X_{1}+X_{10}(t)$ based on 1000 realizations of the Hamiltonian (76) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $L=10$ and $t=50$. The $r$-parameter distribution was fitted using the form (106), yielding the Dyson index $\nu=2.000\pm 0.002$. The solid lines represent (a) the arcsine distribution, (b) the Wigner-Dyson distribution with $\nu=2$, and (c) the Wigner $r$-parameter distribution with $\nu=2$. The corresponding average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% at each edge.
@@
Figure 19: Eigenvalue density, level spacing distribution, and $r$-parameter distribution for the operator $Z_{1}+Z_{10}(t)$ based on 1000 realizations of the Hamiltonian (76) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $L=10$ and $t=50$. The $r$-parameter distribution was fitted using the form (106), yielding the Dyson index $\nu=2.000\pm 0.002$. The solid lines represent (a) the arcsine distribution, (b) the Wigner-Dyson distribution with $\nu=2$, and (c) the Wigner $r$-parameter distribution with $\nu=2$. The corresponding average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% at each edge.
@@
Figure 20: Eigenvalue density, level spacing distribution, and $r$-parameter distribution for the operator $X_{1}+X_{10}(t)$ based on 1000 realizations of the Hamiltonian (76) at the near-integrable point $(h_{x},h_{z},g)=(-1,0.005,0.2)$, with $L=10$ and $t=5\times 10^{4}$. The $r$-parameter distribution was fitted using the form (106), yielding the Dyson index $\nu=2.001\pm 0.002$. The solid lines represent (a) the arcsine distribution, (b) the Wigner-Dyson distribution with $\nu=2$, and (c) the Wigner $r$-parameter distribution with $\nu=2$. The corresponding average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% at each edge.
@@
Figure 21: Eigenvalue density, level spacing distribution, and $r$-parameter distribution for the operator $Z_{1}+Z_{10}(t)$ based on 1000 realizations of the Hamiltonian (76) at the near-integrable point $(h_{x},h_{z},g)=(-1,0.005,0.2)$, with $L=10$ and $t=5\times 10^{4}$. The $r$-parameter distribution was fitted using the form (106), yielding the Dyson index $\nu=2.001\pm 0.002$. The solid lines represent (a) the arcsine distribution, (b) the Wigner-Dyson distribution with $\nu=2$, and (c) the Wigner $r$-parameter distribution with $\nu=2$. The corresponding average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% at each edge.
@@
In finite numerical simulations, free cumulants in the model (76) do not vanish due to finite-size effects. The above results show that while the dynamics drives $Z_{1}$ and $Z_{10}(t)$ towards asymptotic freeness – producing a spectrum for their sum that respects the free probability prediction – the same is not entirely true for $X$-Pauli operators, in which one observes small deviations from the free probability prediction even at the chaotic point. This suggests that the finite size effects for mixed cumulants involving $X$-Pauli operators are larger than the ones involving $Z$-Pauli operator, which is indeed what is observed in numerical simulations. The underlying reason why some operators exhibit larger finite-size effects than others lies in the fact that the Hamiltonian does not treat all operators equally. In the version of the Ising model we consider, for instance, the Hamiltonian takes the schematic form $H=\sum_{i}Z_{i}Z_{i+1}+Z_{i}+X_{i}.$ If we exchange $X\leftrightarrow Z$, we find that cumulants involving $X$ tend to thermalize more effectively (i.e., exhibit smaller residual values) than those involving $Z$.
5.4 Spatial Dependence of $t_{\text{arcsine}}$
In this subsection, we investigate the spatial dependence of $t_{\text{arcsine}}$, the characteristic time scale at which the free probability prediction becomes accurate for operators of the form $Z_{i}(0)+Z_{i+\Delta x}(t)$. Specifically, we analyze how $t_{\text{arcsine}}$ varies with the spatial separation between the operators, defined as $\Delta x$. To account for possible edge effects, we consider both open and closed boundary conditions. The results are shown in Fig. 22.
Since freeness between $Z_{i}(0)$ and $Z_{i+\Delta x}(t)$ indicates that both operators do not commute, having a large commutator, we expect that the arcsine time for distant operators should be larger as compared to operators which are close to each other. Moreover, assuming that the support of the operator $Z_{i+\Delta x}(t)$ grows linearly with time, we expect a linear behavior of $t_{\text{arcsine}}$ with $\Delta x$. In fact, after a non-monotonic behavior for small values of $\Delta x$, we observe that $t_{\text{arcsine}}$ indeed grows linearly $\Delta x$. We expect the delay on the onset of the free probability prediction to be controlled by $\Delta x/v_{F}$, where $v_{F}$ is some velocity scale. From the linear fit we estimate $v_{F}=0.70\pm 0.03$ for $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$ and system size $L=10$. It would be interesting to compare $v_{F}$ with the butterfly velocity extracted from OTOCs. However, a precise comparison requires much larger system sizes and is beyond the scope of this paper.
For both open and closed boundary conditions, the arcsine time exhibits a non-monotonic dependence on the spatial separation for small values of $\Delta x$. Specifically, $t_{a}$ initially decreases with increasing $\Delta x$, reaching a minimum at $|\Delta x|=2$, and subsequently grows linearly, as expected. One possible explanation for this behavior is that the presence of the Ising term $Z_{i}Z_{i+1}$ facilitates the emergence of freeness between a time-evolved operator and its neighboring operators – causing it to fail strongly to commute with them – more readily than with itself at time zero. However, as more distant operators are considered, the expected delay in the onset of the free probability regime emerges, since the local Ising interaction cannot influence distant points instantaneously.
Refer to caption Refer to caption
(a) Open boundary conditions (b) Closed boundary conditions
Figure 22: Arcsine time $t_{\text{arcsine}}$ as a function of spatial separation $\Delta x$, illustrating the emergence of free probability behavior for operators of the form $Z_{i}(0)+Z_{i+\Delta x}(t)$. Results correspond to the chaotic regime with parameters $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$ and system size $L=10$. The arcsine time was extracted via a $\chi^{2}$ analysis over 50 independent realizations of the Hamiltonian (76). A linear fit for $\Delta x\geq 3$ yields $t_{\text{arcsine}}=a+b\,\Delta x$, with $a=3\pm 4$ and $b=1.42\pm 0.07$. Panel (a) shows results with open boundary conditions and $i=1$, while panel (b) uses closed boundary conditions and $i=5$.